Circular No. 6333

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Schleicher, Lowell Observatory, communicates: “Narrow-band pho-

tometry of C/1996 B2 was obtained by R. Millis and myself using the Hall 1.07-m and the 0.79-m telescopes at Lowell Observatory on Feb. 25.5 (r = 1.60 AU) and Mar. 1.5 UT (r = 1.51 AU), respectively. Haser-model production rates for the two dates, respectively, were: log Q(OH) = 28.68 and 28.73, log Q(CN) = 26.08 and 26.15, and log Q(C_{2}) = 26.22 and 26.32. The dust production rates were log (Afρ) = 3.38 and 3.45. Combining these results with those previously obtained on Feb. 9.5 at r = 1.90 AU (cf. IAUC 6311) implies that the OH- and CN-production rates are essentially unchanged over the three-week time span, while C_{2} increased by 50 percent and dust production has increased by ∼ 80 percent. Therefore, while the water-production r dependence is apparently flat, the species primarily responsible for naked-eye visibility currently exhibit a production-rate slope of ∼ −2. This assumes that any short-term variations are not distorting the true r dependence, an assumption consistent with recent IUE observations (see below).

M. Festou, H. Andernach, M. A’Hearn, M. Haken, G. Moreels, and J. Clairemidi report for the International Ultraviolet Explorer (IUE) ToO Team the following observations of C/1996 B2 made at VILSPA on Feb. 19.3 (r = 1.71 AU), 22.2 (1.66), and 27.2 UT (1.57): “The production of both water and dust shows no change over the interval. Using the same Haser model as that of Schleicher and Millis (see above), log Q(OH) = 28.68, 28.68, and 28.64, in remarkably good agreement with their results. For a ‘best’ vectorial model, the corresponding values of Q would be a bit more than 2 times larger. The dust production at 300–304 nm can be characterized by log Afρ = 3.17, 3.17, and 3.18 (cf. A’Hearn et al. 1984, AJ 89, 579). These values are not directly comparable with the values found by Schleicher and Millis above, since they are measured at a different wavelength. Taken together, they suggest that the dust has changed from neutral in color to reddish. CS is also approximately constant over the interval. This complete lack of brightening as the comet approaches the sun is just the behavior that might have been expected for a dynamically new comet, although the latest orbital solutions (Marsden, IAUC 6329) suggest that the comet is not dynamically new.”

1996 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green